Problem 3 Solve the initial value problems using Laplace Transforms (a) y + 8y = t2 y(0) = -1 (b) y – 2y – 3y = e4t y(0) =
Answer #1

Consider y+8y=t y (0) = -) Take Taplas transform on both side £1 y 3 + hrdy3 = 293 sy(s)- 4/0) + Py(s) = 2 sy(s) +1 +84(s) a. + $t inscence (if feats, + +*+ 997 e pt to so :: ylte + + - 257 + DS 256 3 y- Qy - 34 ze4+ y(ol=1, 7! (0)=-) Taking lapla3465) . (s. 14) (Ŝ-25-3) 3-75+12 +1 ? :-4) (5-3) (5-4) (5-3) (5+1) t ! (5-4) (5-3) (5+1) (5-4) (5-3) (st!! 5+. (54) (5-3)(5+1